bottled water: from 1990 to 1999, the per person consumption B of bottled water (in gallons) and the population P of the united states (in thousands) can be modeled by
B = 0.0977t2 + 0.186t + 7.86 and
P = 3226t+ 250,359
where t is the number of years since 1990. write a model for the total consumption C of bottled water (in thousands of gallons). what is the total consumption of bottled water in 1998?

Answer by stanbon(75887)

bottled water: from 1990 to 1999, the per person consumption B of bottled water (in gallons) and the population P of the united states (in thousands) can be modeled by
B = 0.0977t^2 + 0.186t + 7.86 and
P = 3226t+ 250,359
where t is the number of years since 1990. write a model for the total consumption C of bottled water (in thousands of gallons). what is the total consumption of bottled water in 1998?
-----------------
Total consumption = B*P
---
I wouldn't try to simplify that product as it is too messy.
----
Total(t) = (0.0977t^2 + 0.186t + 7.86)*(3226t+ 250,359)
----
In 1998 t = 8. So Find Total(8) and use your calculator to get:
T(8) = 15.6008*276167 = 4308426.134
=========================================
Cheers,
Stan H.
)